Prev Next

SW&IT Meeting Minutes 2018-10-01

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • EM interview

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • Luci H/O doc (IS)
  • Luci telemetry
    • web-io cold heads state for vibration (IS)
    • luci telemetry in git (DONE)
  • vm backup

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary Week
OAC
AC (3)
Petr (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (4)
Petr (2)
2
Weatherstation
AC (2)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
ALH Igor(1/2)
Matthieu (2)
1
INDI Igor Matthieu 3
GCS
Petr (3)
Matthieu (2)
1
Telemetry
Petr (3)
Matthieu (2)
2
OVMS
Petr (3)
Matthieu (2)
3
Status Server Stephen (2)
Matthieu (2)
2
Telemetry cleanup Stephen (1) Petr 3
LBC
Yang (1/2)
Igor(1/2) 2
MODS
Yang (1)
Igor 2
Telem-VIZ Yang Stephen (1) 3
particle monitors AC Petr 4

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • 4Runner IP
  • FOSWiki transition
  • Magic laptop clone/vm
  • SOUL/FLAO Oct18 in lab then telescope
  • LMS tcl/tk?
  • INDI / AO

This week in github

Team reports

YZ

  • Met with Kellee, Igor, and Olga on LBC and MODS handover.
    • LBC: major components and structures are clear now, many more wiki documents to go through (over the time).
    • MODS: some knowledge of current use and situation, still many documents to go through.
  • Finished a new version of fitting tool for collimation (supports many more functionality).
  • TODO or continue doing:
    • Working on telemetry search tool (this may be a good opportunity to combine with the work of revamping the current telemetry web interface).
    • Some background info (e.g., who uses telemetry search the most and what they'd like to do, etc.). Also software architecture design may need to be discussed as well.
    • May schedule another meeting with Kellee sometime to discuss other stuffs (MODS, current telemetry web interface etc.).
    • Continue on collimation documentation.
    • Meet with Al Conrad to discuss more details of the AO precheck project.

PK

  • some edits to GCS and Telemetry, still way to finish
  • with AC - OAC/AGW4 problems - guide probe help, with:
    • readline integration of pmacsend
    • telemetry support basis, so we understand a whole more about AOC (RPC calls, AOCONTROL threads)
    • fixed 32/64 bit aocontrol build
    • code inmpovements (const char* instead of char* where approprite - adds readibility)

AC

Last week:
  • ARGOS internal meeting
    • action item: transition Arcetri wiki to FOSwiki by 10/11
  • OAC/AGW4 (see PK)
  • HO tech talk review with KS

This week:
  • Continue OAC/AGW4
  • Continue HO tech talk reviews
  • Review ARGOS ITs: 7162, 5453, 5714
  • Wiki transition

MB

  • management OC, WPC
  • TMT ESW review
  • cisrv01 recovery
  • IA2 next meeting Oct 18
  • ALTA telemetry
  • Safety trainingfor mgr

next:
  • EngMgr interviews Thu, 3rd candidates
  • IT to github issues (overdue), IT reassignments
  • OVMS+TT AO network

SH

Last week:
  • foswiki
  • single-mode for LBTI
  • DHCP issues
  • Jenny TCL
  • cisrv01 reinstall
  • power outage on Tuesday

Next:
  • routing for Petr -- mountain IPs in downtown lab
  • Docker repo
  • LBTI? ports will be needed
  • MGIO? need to talk to Eric

KS

Last week

  • OAC
    • IT from PK about logging/telemetry. Turned off debug and rebuilt the mountain version. Telemetry was fine - he was looking in Tucson.
    • noticed the AGW4 alarms were triggered for voltages - this was a bug in the tcs.db - fixed and restarted the IOCs with Igor.
    • went over the UMAC telemetry required with AC.
  • Pulled updated leap seconds file and passed on to the team
  • Status Server
    • Reviewed SH's Status Server questions and wrote up response.
    • Checked in to SVN the INDI AO drivers and indi-fcgi source.
  • LBC/MODS discussion with Yang and Igor - active optics, build procedure, MODS machines, etc.
  • Reviewed Petr's first cut at handover techtalks from his main pages for OVMS, GCS, telemetry.
  • Discussion with Andrew about how to run DIMM and where the operating procedures are on the wiki since Matthieu tried to run it and couldn't find the relevant info, so it was failing to come up. Reviewed Andrew's first cut at a DIMM techtalk. Discussed outstanding updates and build procedures.
  • Reviewed all my DIMM hardcopy notes, cannot find the updated DIMM Operating Principles document that I have a hardcopy of. The one in vault is older.

This week
  • Update the DIMM operator's manual with auto-acquisition details.
  • Continue going through all my application notes, binders, wiki pages, email to pass on relevant info.
  • Vacation on Thursday.

DE

Last week:
  • Finish setting up VMWare tools on VMs for Veeam usage
    • One outstanding due to complications with reboot, causing warning messages but we are getting backups.
  • Unarchive tuc-queueserv1 at Igor's request
    • He's able to log in
    • Re-added it to DNS
  • Peter laptop setup
    • Dell still hasn't fixed the laptop/dock driver issues out of the box, something to keep in mind for future orders
  • Matthieu laptop repair setup
  • Set up network mount for Vault backups to be written to Synology
    • Waiting on Ketiv to change backup location
  • Set up offsite replication of VM backups and archive data to Google Drive for better disaster recovery
This week:
  • Test restores of selected VMs to make sure all recent changes are captured
  • Reminder: Fire alarm tests Monday and Tuesday 0700-1500, state only one will be audible.
  • Time got away from us re: quick meeting about server room power; what day is good for try 2?

IS

Last week:

Mostly LUCI stuff, some PIT
  • LUCI:
    • General support issues:
      • how commands are sent to mirror tilting device
      • how commands are sent and how often to filter device (on going)
    • Development:
      • Working on getting all the LMS information on the FITS header, at the time the FITS gets generated by GEIRS and reading it back on the RTD, ETA 2 to 3 days
  • PIT
    • Test and release
  • Handover instruments
    • LBC and MODS meeting with Kellee, Yang, Olga
This Week:

Mostly LUCI work and handover instruments
  • LUCI
    • Finishing RTD work
    • Looking at how the filter commands are generated
    • Possibly try to look at how simplify dev environment
    • Journalizer innerworkings. Create a simple tool that generates and reads Journalizer events.
  • OT/PIT
    • Work with Michelle on OT/PIT unfinished work.
  • Handover
    • INDI
    • LBC/MODS
Topic revision: r18 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback